Time Frequency Brown Bag Seminar Wednesday May 20, 1998 12:00 noon EQuad E415 Speaker: James Wei Dean, School of Engineering And Applied Science Princeton University Title: Mixing: Measures of the Structure of Uniformity Abstract:

There is a great deal of literature on the process of mixing, but almost silent on outcome as assessment. Think of a Cuisinart full of lettuce, then a small carrot is added, and we have data of red color as c(x, y, z, t). When is the mixture sufficiently uniform? I will propose a number of measures, at various scales. What is the appropriate length scale? I will illustrate with some examples from the literature.